# CHOICE: Coordinated Human-Object Interaction in Cluttered Environments for Pick-and-Place Actions

### v1

@@ 9. Discussion and Future Works
In this work, we demonstrate the effectiveness of our implicit neural trajectory planner in generating realistic pick-and-place guidance with strong generalization across various layouts. Combined with a robust goal-driven controller that simultaneously estimates the tracking state in the frequency domain, our system drives the character to perform complex interactions guided by goals of three key joints. This capability is further enhanced by our bimanual scheduler. Together, these techniques are hierarchically integrated to form the first systematic method for long-term interaction synthesis in cluttered environments. Despite these advances, our pipeline has certain limitations as follows.
@@
The characters are currently controlled by mouse-clicking; which is intuitive but requires some effort as the viewpoint also may need to be managed by the user. It will be interesting to look into applying LLMs to control the character by using language as an interface.
Additionally, our full-body trajectory planning for navigating towards the goal is currently based on a 2D path planner. Thus, the character cannot avoid obstacles that require the body to bend down, tilt the torso to the side or walk over. One possibility is to incorporate sensors that evaluate the distances to 3D objects for path planning and to use such sensors for motion synthesis during runtime (Holden et al., 2017; Starke et al., 2019).

### v2

@@ 9. Discussion and Future Works
In this work, we demonstrate the effectiveness of our implicit neural trajectory planner in generating realistic pick-and-place guidance with strong generalization across various layouts. Combined with a robust goal-driven controller that simultaneously estimates the tracking state in the frequency domain, our system drives the character to perform complex interactions guided by the goals of three key joints. This capability is further enhanced by our bimanual scheduler. Together, these techniques are hierarchically integrated to form the first systematic method for long-term interaction synthesis in cluttered environments. The proposed framework can be extended in several ways to tackle even more complex scenarios.
@@
The characters are currently controlled by mouse-clicking, which is intuitive but requires some effort, as the viewpoint also may need to be managed by the user. It will be interesting to look into applying LLMs to control the character by using language as an interface.
Limitations
Despite the strengths exhibited in our evaluations, our pipeline has certain limitations as follows.
Navigation-Induced Collisions
Our full-body navigation towards the goal is currently based on a 2D path planner and navigation matching. Thus, the character cannot well avoid obstacles that require the body to bend down, tilt the torso to the side, or walk over. As a result, the reaching motion generated through tracking a sequence of keyjoint goals is not guaranteed to be collision-free between the goal waypoints. It is difficult to resolve collisions in an online manner, as this can degrade the motion coordination and smoothness (see Fig. 13(a) and the supplementary video). One possibility is to incorporate sensors that evaluate the distances to 3D objects for the 2D path planning and to use such sensors for motion synthesis during runtime (Holden et al., 2017; Starke et al., 2019).
Generalization Gaps to Unseen Furniture
Performing collision-free and coordinated human interaction in arbitrary cluttered environments presents significant challenges. Our current system exhibits limitations in generalization to unseen furniture that is significantly different from that in our dataset (see Fig. 13). For full-body motion, shifting the 3-keyjoint goal transformations for collision-resolving (described in § 6.3) can lead to uncoordinated motion. E.g., in Fig. 13(b), the increased hand-hip distance imposed by the oven door results in an unnatural full-body motion with foot sliding. Additionally, on the manipulating hand, collisions sometimes occurred due to the failed trajectory planning, when it cannot effectively guide the manipulating hand to pass through surrounding furniture obstacles. For instance, in Fig. 13(c), an ideal trajectory should guide the hand to reach into and depart from the box’s opening while avoiding the wineglass.
However, these issues can be alleviated through computing an optimal goal-reaching pose adaptive to diverse scenes, by leveraging prior knowledge extracted from a vast video dataset of human interactions. For example, instead of approaching from the front, as shown in Fig. 13(b,c), the character could potentially reach the target more closely from less obstructed sides, thereby mitigating the conflict between humanoid collision avoidance and realistic grasping. To achieve this, we could first learn a probabilistic 2D pose prior from extensive, unlabeled interaction videos, which acts as a realism evaluator. A plausible goal pose could then be found by programming in the canonical space, with the objective of maximizing the score aggregated from this evaluator across multiple 2D projections, subject to the test-time collision constraints.
Acknowledgements.
This project was partially funded by Meta. It was also partially funded by the Research Grants Council of Hong Kong (Ref: 17210222), and by the Innovation and Technology Commission of the HKSAR Government under the ITSP-Platform grant (Ref: ITS/335/23FP) and the InnoHK initiative (TransGP project). Part of the research was conducted in the JC STEM Lab of Robotics for Soft Materials, funded by The Hong Kong Jockey Club Charities Trust.